OpenStudy (anonymous):

What Is (-m)^-3n If m= 2 and n= -24? A. 3 B. -3 C. 4 D. -4

OpenStudy (ash2326):

(-m)^(-3)* n (-2)^(-3) *(-24) a^(-b)=1/a^(b) so (-2)^(-3)=1/(-2)^3 1/(-2)^3 * -24 -24/ -8= 3 option A

OpenStudy (anonymous):

Thank You (: